This one should be indented, hhh hhhhhh hh hhhhhhh hhhhhh hhhh hhhhhh hhhhhhh hhhhhhhh hhhhhhh hhhhh

1. This is the first item, and spans only one paragraph. It has no left indentation, as you can see.

2. This is the second item, and it spans two paragraphs. This is the first paragraph and has no indentation.

qqqqqq This is the second paragraph, and is indented. But it has no numbering, because it belongs to the same enumeration item, the second one.

3. last item

Will this one be indented ? I guess no. hhh hhhhhh hh hhhhhhh hhhhhh hhhh hhhhhh hhhhhhh hhhhhhhh

This one should be indented, hhh hhhhhh hh hhhhhhh hhhhhh hhhh hhhhhh hhhhhhh hhhhhhhh hhhhhhh hhhhh